what is the rectangular equivalent of a parametric curve....x=t/5-1 and y=t-21
Answer by stanbon(75887)

what is the rectangular equivalent of a parametric curve....
x= (t/5)-1 and y=t-21
--------------------
Solve each equation for "t":
t/5 = x+1
t = 5x+5
---------------
t = y+21
------------------
Equation:
5x+5 = y+21
y = 5x-16
=================
Cheers,
Stan H